OpenStudy (anonymous):

The world population in the year 2011 was approximately 7 x 109 people. If the population doubles in 100 years, how many people will be on the earth in the year 2111? Write your answer in scientific notation.

OpenStudy (kropot72):

Population in 2011 = 7 * 10^9 Population in 2111 = 2 * 7 * 10^9 Now you can work out the answer.